If x tan 30°+ y cot 60° = 0 and 2x-y tan 45°=1 then find the value of x and y.Thank you for solving it but I want it to be solved simultaneously

Grade:11

1 Answers

Arun
6 years ago
Dear Shraddha
 
x / sqrt(3) + y / sqrt(3) = 0
 
hence
x = – y
 
now
2x – y =1
now
2x – ( – x) = 1
2x + x= 1
3x = 1
x = 1/3
 
now y = – 1/3
 
Regards
Arun (askIITians forum expert)